A point M is lying 25 mm above the H.P and 15 mm behind the vertical plan. Another point N is lying 30 mm below the H.P and 20 mm in front of the V.P. The distance between their projectors is 40 mm. Draw their projections and find the angle which a line joining their top views extends with xy reference line.
Answer #1


* Perojections a M ī 25mm M X o OG neve heus ac Z zk k here on 부 joining Angle their which line TOP views is Q. Let ve us fi 15 20 40-22 2 15x2 = 800 - 20X2 i. 35 X2 2800 ره 22.86 mm x = 40 x2 =40-22.56 4 = 17.14 mm cuve si E Angle o - tano ho o tanc